Solved by lagrange method

y^2 p^2+x^2q^2=x^2y^2z^2


1
Expert's answer
2021-01-13T19:00:05-0500

"F(x,y,z,p,q)=y^2p^2+x^2q^2-x^2y^2z^2=0"

"\frac{dx}{F_p}=\frac{dy}{F_q}=\frac{dz}{pF_p+qF_q}=-\frac{dp}{F_x+pF_z}=-\frac{dq}{F_y+qF_z}"


"\frac{dx}{2py^2}=\frac{dy}{2qx^2}=\frac{dz}{2p^2y^2+2q^2x^2}=-\frac{dp}{2xq^2-2xy^2z^2-2pzx^2y^2}=-\frac{dq}{2yp^2-2yx^2z^2-2qzx^2y^2}"


"\frac{pdx+qdy-dz}{2p^2y^2+2q^2x^2-2p^2y^2-2q^2x^2}=\frac{pdx+qdy-dz}{0}=-\frac{dp}{2xq^2-2xy^2z^2-2pzx^2y^2}=-\frac{dq}{2yp^2-2yx^2z^2-2qzx^2y^2}"


"\frac{dx}{2py^2}=\frac{dy}{2qx^2}=\frac{pdx+qdy-dz}{0}"

"p=0, q\neq0"

Substitute this result into the initial equation, and we get:

Then:

"x^2q^2=x^2y^2z^2"

"q=yz"

"\frac{y^2}{2}=lnz+c_1"


"p\neq0, q=0"

Then:

"p=xz"

"\frac{x^2}{2}=lnz+c_2"


The general solution:

"F(c_1,c_2)=F(\frac{y^2}{2}-lnz, \frac{x^2}{2}-lnz)"